Two identical balls each moving with a speed v at right angle to each other collide perfectly inelastically. Their speed after collision is?

Grade:11

1 Answers

Khimraj
5 years ago
let spped after collision is x then
apply conservation of momentum
mv\hat{i} + mv\hat{j} = 2mx
so x = \sqrt{2}v/2 = v/\sqrt{2}
Hope it clears.